OpenStudy (anonymous):

Solve the equation. (Enter all answers including repetitions. If there is no solution, enter NO SOLUTION.) x^4 - 32x^2 + 256 = 0 x= (smallest value) x= x= x= (largest value)

hero (hero):

x^4 - 16x^2-16x^2 + 256 =0 x^2(x^2 -16)-16(x^2-16) = 0 (x^2-16)(x^2-16) = 0 I think you can solve for x from there

OpenStudy (anonymous):

how?

hero (hero):

x^2 - 16 = 0 x^2 = 16 x = +/- 4

hero (hero):

there's four values x = -4 x = -4 x = 4 x = 4